A printing company has to print a document that conatins a total of 48,000 words. They have only two sizes of typeset to work with: large, 1800 words per page, and small, 2400 words per page. The customer wants to have as many pages with large type set as possible. The entire document cannot be more than 21 pages total. How many pages of large type can they print?
Answer by stanbon(48558)

Let number of large type pages be "x"
Then number of small type pages is "21-x"
# of words on large type pages is 1800x
# of words on small type pages is 2400(21-x)
EQUATION:
# of words + # of words = 48000
1800x+2400(21-x)=48000
1800x+59400-2400x=48000
-600x=-2400
x=4 (# of pages with large type)